OpenStudy (anonymous):

Completely factor the following expression81y^2-4z^2

OpenStudy (anonymous):

difference of two squares take sqrt of each of the terms use a^a + b^2 = (a + b)(a - b) as a template

OpenStudy (anonymous):

Get the square root of \[ \sqrt{81y^2} \] = 9y this is A Get the square root of \[ \sqrt{4z^2} \] = 2z this is B Now just plug the numbers into (a + b)(a-b) (9y+2z)(9y - 2z)